### v2

@@ Appendix B Omitted proofs from Section 5
Assume then that we know the correspondence $\hat{\mathrm{DS}}_{t}$ is nonempty valued for $t\in\{1,\dots,T-1\}$. As it is clear from the main text, to establish the main result, showing that for any $E^{T}$, the period-1 conjectures $\{\varphi_{\hat{\mathrm{DS}},1}(\cdot,k):k\in A_{1}\cup B_{1}\}$ are nonempty and satisfy consistency is enough.
@@
We observe that for all $k\in A_{1}\cup B_{1}$, the following hold. First, $\varphi_{n+1}(\emptyset,k)\subseteq\varphi_{n}(\emptyset,k)$ because in each step we are increasing the continuation values of period-1 agents other than $k$. Formally, by definition we have that for all $k\in A_{1}\cup B_{1}$, $\varphi_{1}(\emptyset,k)\subseteq\varphi_{0}(\emptyset,k)$. Suppose we have shown that $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$ for all $k^{\prime}\in A_{1}\cup B_{1}$; we proceed to show that $\varphi_{n+1}\subseteq\varphi_{n}$. Fix $k\in A_{1}\cup B_{1}$ and $\overline{m}\in\varphi_{n+1}(\emptyset,k)$. We only need to show that if $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n}}^{\overline{m}_{1}})$, then $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}})$. To this end, let $k^{\prime}$ be such that $\overline{m}_{1}(k^{\prime})\neq k^{\prime}$. By construction, $k^{\prime}$ prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $k^{\prime}$ also prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $\overline{m}\in\varphi_{n}(\emptyset,k)$. Second, because everything is finite $\varphi_{\infty}(\emptyset,k)=\lim_{n\rightarrow\infty}\varphi_{n}(\emptyset,k)$ is well-defined. Third, because $\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}\left(E_{2}^{T}(\overline{m}_{1})\right)\}\subseteq\varphi_{n}(\emptyset,k)$ for all $n$ and is nonempty, then $\varphi_{\infty}(\emptyset,k)$ is nonempty.
We claim that
@@
To see that Equation B.2 holds, note the following:
@@
    Because the sets $\varphi_{n}(\emptyset,k)$ are finite and so is the set of period-1 agents, $\overline{N}<\infty$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}=\varphi_{\overline{N}+1}$. Toward a contradiction, suppose that for all $N\in\mathbb{N}$ $k_{N}\in A_{1}\cup B_{1}$ exists such that $\varphi_{N}(\emptyset,k_{N})\neq\varphi_{N+1}(\emptyset,k_{N})$. Since the set $A_{1}\cup B_{1}$ is finite, we can find an agent $k$ and a subsequence $(n^{k})$ such that $\varphi_{n^{k}}(\emptyset,k)\neq\varphi_{n^{k}+1}(\emptyset,k)$. Thus, $\varphi_{n^{k}+1}(\emptyset,k)\subsetneq\varphi_{n^{k}}(\emptyset,k)$. Since $\varphi_{n^{k}}(\emptyset,k)$ is finite, this means that eventually $\varphi_{n^{k}}(\emptyset,k)=\emptyset$, a contradiction. From now on, let $\overline{N}$ denote the smallest $n$ such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)\equiv\varphi_{\infty}(\emptyset,k)$.
@@
    Suppose $k\in A_{1}\cup B_{1}$ exists such that $\overline{m}\in\varphi_{\infty}(\emptyset,k)$, but $\overline{m}$ is not an element of the set on the right-hand side of Equation B.2. Then, without loss of generality, $a\in A_{1}\cup B_{1}\setminus\{k\}$ exists such that $\overline{m}_{1}(a)\neq a$ and $\min_{\tilde{m}\in\varphi_{\infty}(\emptyset,a)}U_{1}(a,\tilde{m})>u(a,\overline{m}_{1}(a)).$ This implies that $\overline{m}\notin\varphi_{\overline{N}+2}(\emptyset,k)$, contradicting that $\varphi_{\overline{N}+1}(\emptyset,k)=\varphi_{\infty}(\emptyset,k)$.
@@
We conclude that for all $k\in A_{1}\cup B_{1}$, Equation B.2 holds and hence $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$. Thus, for all $k\in A_{1}\cup B_{1}$, $\varphi_{\hat{\mathrm{DS}}}(\emptyset,k)\neq\emptyset$.
@@
We now show that $\varphi_{\hat{\mathrm{DS}},1}$ satisfy CC. To this end, construct a candidate $m^{\star}$ as described in Section 4 using $\varphi_{\hat{\mathrm{DS}},1}$ and the correspondence $\hat{\mathrm{DS}}_{T-1}$, which by the inductive hypothesis is nonempty. Fix a period-1 agent, $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$. Note that $m^{\star}$ satisfies the following: First, the continuation matching $(m^{\star}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}(E^{T}_{2}(m^{\star}_{1}))$. Second, for all $k^{\prime}\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k^{\prime})\neq k^{\prime}$, we have that $k^{\prime}$ is matched to someone they prefer to the worst matching in $\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k^{\prime})$. Finally, note that $m^{\star}_{1}$ has no blocking pairs $(a,b)$ such that $m^{\star}_{1}(a)\neq a,m^{\star}_{1}(b)\neq b$. Thus, $m^{\star}_{1}\in\mathrm{S}(E_{1,\varphi_{\hat{\mathrm{DS}}}}^{m^{\star}_{1}})$. Hence, $m^{\star}\in\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k)$ so that CC holds.
@@
Theorem 4.1 then implies that $m^{\star}\in\hat{\mathrm{DS}}_{T}(E^{T})$, which completes the inductive step. ∎

### v3

@@ Appendix B Omitted proofs from Section 5
Assume then we know the correspondence $\hat{\mathrm{DS}}_{t}$ is nonempty valued for $t\in\{1,\dots,T-1\}$. As is clear from the main text, to establish the main result, showing that for any $E^{T}$, the period-1 conjectures $\{\varphi_{\hat{\mathrm{DS}},1}(\cdot,k):k\in A_{1}\cup B_{1}\}$ are nonempty and satisfy consistency is enough.
@@
We observe that for all $k\in A_{1}\cup B_{1}$, the following hold. First, $\varphi_{n+1}(\emptyset,k)\subseteq\varphi_{n}(\emptyset,k)$ because in each step, we are increasing the continuation values of period-1 agents other than $k$. Formally, by definition, we have that for all $k\in A_{1}\cup B_{1}$, $\varphi_{1}(\emptyset,k)\subseteq\varphi_{0}(\emptyset,k)$. Suppose we have shown $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$ for all $k^{\prime}\in A_{1}\cup B_{1}$; we proceed to show $\varphi_{n+1}\subseteq\varphi_{n}$. Fix $k\in A_{1}\cup B_{1}$ and $\overline{m}\in\varphi_{n+1}(\emptyset,k)$. We only need to show that if $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n}}^{\overline{m}_{1}})$, $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}})$. To this end, let $k^{\prime}$ be such that $\overline{m}_{1}(k^{\prime})\neq k^{\prime}$. By construction, $k^{\prime}$ prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $k^{\prime}$ also prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $\overline{m}\in\varphi_{n}(\emptyset,k)$. Second, because everything is finite, $\varphi_{\infty}(\emptyset,k)=\lim_{n\rightarrow\infty}\varphi_{n}(\emptyset,k)$ is well-defined. Third, because $\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}\left(E_{2}^{T}(\overline{m}_{1})\right)\}\subseteq\varphi_{n}(\emptyset,k)$ for all $n$ and is nonempty, $\varphi_{\infty}(\emptyset,k)$ is nonempty.
We claim
@@
To see Equation B.2 holds, note the following:
@@
    Because the sets $\varphi_{n}(\emptyset,k)$ are finite, as is the set of period-1 agents, $\overline{N}<\infty$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}=\varphi_{\overline{N}+1}$. Toward a contradiction, suppose that for all $N\in\mathbb{N}$ $k_{N}\in A_{1}\cup B_{1}$ exists such that $\varphi_{N}(\emptyset,k_{N})\neq\varphi_{N+1}(\emptyset,k_{N})$. Because the set $A_{1}\cup B_{1}$ is finite, we can find an agent $k$ and a subsequence $(n^{k})$ such that $\varphi_{n^{k}}(\emptyset,k)\neq\varphi_{n^{k}+1}(\emptyset,k)$. Thus, $\varphi_{n^{k}+1}(\emptyset,k)\subsetneq\varphi_{n^{k}}(\emptyset,k)$. Since $\varphi_{n^{k}}(\emptyset,k)$ is finite, this means that eventually, $\varphi_{n^{k}}(\emptyset,k)=\emptyset$, a contradiction. Hereafter, let $\overline{N}$ denote the smallest $n$ such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)\equiv\varphi_{\infty}(\emptyset,k)$.
@@
    Suppose $k\in A_{1}\cup B_{1}$ exists such that $\overline{m}\in\varphi_{\infty}(\emptyset,k)$, but $\overline{m}$ is not an element of the set on the right-hand side of Equation B.2. Then, without loss of generality, $a\in A_{1}\cup B_{1}\setminus\{k\}$ exists such that $\overline{m}_{1}(a)\neq a$ and $\min_{\widetilde{m}\in\varphi_{\infty}(\emptyset,a)}U_{1}(a,\widetilde{m})>u(a,\overline{m}_{1}(a)).$ This implies $\overline{m}\notin\varphi_{\overline{N}+2}(\emptyset,k)$, contradicting that $\varphi_{\overline{N}+1}(\emptyset,k)=\varphi_{\infty}(\emptyset,k)$.
@@
We conclude that for all $k\in A_{1}\cup B_{1}$, Equation B.2 holds, and hence, $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$. Thus, for all $k\in A_{1}\cup B_{1}$, $\varphi_{\hat{\mathrm{DS}}}(\emptyset,k)\neq\emptyset$.
@@
We now show $\varphi_{\hat{\mathrm{DS}},1}$ satisfy CC. To this end, construct a candidate $m^{\star}$ as described in Section 4 using $\varphi_{\hat{\mathrm{DS}},1}$ and the correspondence $\hat{\mathrm{DS}}_{T-1}$, which by the inductive hypothesis is nonempty. Fix a period-1 agent, $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$. Note $m^{\star}$ satisfies the following: First, the continuation matching $(m^{\star}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}(E^{T}_{2}(m^{\star}_{1}))$. Second, for all $k^{\prime}\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k^{\prime})\neq k^{\prime}$, we have that $k^{\prime}$ is matched to someone they prefer to the worst matching in $\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k^{\prime})$. Finally, note $m^{\star}_{1}$ has no blocking pairs $(a,b)$ such that $m^{\star}_{1}(a)\neq a,m^{\star}_{1}(b)\neq b$. Thus, $m^{\star}_{1}\in\mathrm{S}(E_{1,\varphi_{\hat{\mathrm{DS}}}}^{m^{\star}_{1}})$. Hence, $m^{\star}\in\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k)$ so that CC holds.
@@
Theorem 4.1 then implies $m^{\star}\in\hat{\mathrm{DS}}_{T}(E^{T})$. Because the proof only used that $m^{\star}$ is a candidate matching, this completes the inductive step. ∎
Proof of 5.2.
The proof proceeds by complete double induction on the length of the economy, $T$, and the number of agents arriving in period 1, $n$. Let $\mathrm{P}(T,n)$ denote the following inductive statement:
$\mathrm{P}(T,n)$: For all economies of length $T$ with $n$ agents in period $1$, $\mathrm{SDS}_{T}\subseteq\mathrm{DS}_{T}$.
That $\mathrm{P}(1,n)=1$ for all $n\in\mathbb{N}$ follows from noting that for all length-1 economies, $E_{1}$, $\mathrm{SDS}_{1}(E_{1})=\mathrm{DS}_{1}(E_{1})=\mathrm{S}_{1}(E_{1})$. Similarly, $\mathrm{P}(2,1)=1$ because in a length-2 economy $E^{2}$ in which only one agent arrives in $t=1$, the only candidates for (sophisticated) dynamically stable matchings are matchings $(m_{1},m_{2})$ whose continuation matching is stable for the length-1 economy $(A_{1}\cup A_{2},B_{1}\cup B_{2})$. That is, for any length-2 economy in which only one agent arrives in period 1, the sets of dynamically stable and sophisticated dynamically stable matchings coincide.
The rest of the proof consists of the following two inductive steps. First, if $T\geq 1$ and $n\geq 1$ exist such that, for all $n^{\prime}\leq n$, $\mathrm{P}(T,n^{\prime})=1$, then $\mathrm{P}(T,n+1)=1$. This first step then proves that $\mathrm{P}(T,\cdot)=1$. Second, if $T$ exists such that, for all $T^{\prime}<T$, $\mathrm{P}(T^{\prime},\cdot)=1$, $\mathrm{P}(T+1,1)=1$. Together, these steps prove that the inductive statement is true for all $T\in\mathbb{N}$ and $n\in\mathbb{N}$.Footnote 20Footnote 20Footnote 20Indeed, $\mathrm{P}(1,\cdot)=\mathrm{P}(2,1)=1$ by our initial steps. Moreover, $\mathrm{P}(2,1)=1$ and the first inductive step shows that $\mathrm{P}(2,\cdot)=1$. The second inductive step then proves $\mathrm{P}(3,1)=1$. Applying the first and second inductive steps sequentially shows $\mathrm{P}(3,\cdot)=1$, and hence, $\mathrm{P}(4,1)=1$, respectively. Iterating this process shows $P(T,n)=1$ for all $T$ and $n$.
Fix $T\geq 2$ and $n\geq 1$. Fix an economy of length $T$ with $n$ agents in $t=1$, and denote it by $E^{T}$. By the inductive hypothesis, suppose the following two properties hold. First, for all $T^{\prime}<T$ and all $n^{\prime}\in\mathbb{N}$, $\mathrm{SDS}_{T^{\prime}}(E^{T^{\prime}})\subseteq\mathrm{DS}_{T^{\prime}}(E^{T^{\prime}})$ for all economies with $n^{\prime}$ agents in $t=1$. By the discussion above, this implies $P(T,1)=1$, and hence, we can take $n\geq 2$ without loss of generality. Second, for all economies of length $T$ with $n^{\prime}<n$ agents in $t=1$, $E^{\prime^{T}}$, $\mathrm{SDS}_{T}(E^{\prime^{T}})\subseteq\mathrm{DS}_{T}(E^{\prime^{T}})$. We show $\mathrm{SDS}_{T}(E^{T})\subseteq\mathrm{DS}_{T}(E^{T})$.
It suffices to show that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\mathrm{SDS},1}(E^{T},k)\subseteq\varphi_{\mathrm{DS},1}(E^{T},k)$, where we now make apparent that the conjectures in dynamic stability depend on the economy. To do so, we show that for all $l\in\mathbb{N}_{0}$, $F_{l}(E^{T},k)\subseteq\varphi_{\mathrm{DS},1}(E^{T},k)$.
Consider first $l=0$. Recall that
$F_{0}(E^{T},k)$ =$=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k;E^{T}):(\overline{m}_{1}\setminus k,(\overline{m}_{s})_{s=2}^{T})\in\mathrm{SDS}_{T}(E^{T}_{\setminus\left\{k\right\}})\right\}$
$\subseteq\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k;E^{T}):(\overline{m}_{1}\setminus k,(\overline{m}_{s})_{s=2}^{T})\in\mathrm{DS}_{T}(E^{T}_{\setminus\left\{k\right\}})\right\}$
$\subseteq\varphi_{\mathrm{DS},1}(E^{T},k),$ (B.3)
where the first inclusion follows by assumption, and the second inclusion, from observing that dynamic stability imposes more conditions than the conjectures of agent $k$ under dynamic stability on the period-$1$ matching.
The above inclusion implies
$M^{\star}(\emptyset,F_{0},\mathrm{SDS}_{T-1})$ =$=\{\overline{m}\in\pazocal{M}_{T}(E^{T}):\text{ (i) }\overline{m}_{1}\in\mathrm{S}(E_{1,F_{0}})\text{ (ii) }(\overline{m}_{s})_{s=2}^{T}|_{E^{T}_{2}(\overline{m}_{1})}\in\mathrm{SDS}_{T-1}(E^{T}_{2}(\overline{m}_{1}))\}$
$\subseteq\{\overline{m}\in\pazocal{M}_{T}(E^{T}):\text{ (i) }\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi^{DS}})\text{ (ii) }(\overline{m}_{s})_{s=2}^{T}|_{E^{T}_{2}(\overline{m}_{1})}\in\mathrm{DS}_{T-1}(E^{T}_{2}(\overline{m}_{1}))\}$
$\subseteq\mathrm{DS}_{T}(E^{T}),$ (B.4)
where the first inclusion follows from $F_{0}(E^{T},k)\subseteq\varphi_{\mathrm{DS},1}(E^{T},k)$ for all $k\in A_{1}\cup B_{1}$ and $\mathrm{SDS}_{T-1}\subseteq\mathrm{DS}_{T-1}$, and the second inclusion follows from noting the second line describes the candidate matchings for dynamic stability. Thus,
$F_{1}(E^{T},k)=F_{0}(E^{T},k)\cup\{\overline{m}\in M^{\star}(\emptyset,F_{0},\mathrm{SDS}_{T-1}):\overline{m}_{1}(k)=k\}\subseteq\varphi_{\mathrm{DS},1}(E^{T},k),$
where the inclusion follows from Equations B and B. Repeating the argument in Appendix B for $l\geq 1$ delivers that $\varphi_{\mathrm{SDS},1}(E^{T},k)\subseteq\varphi_{\mathrm{DS},1}(E^{T},k)$ and hence that $\mathrm{SDS}_{T}(E^{T})\subseteq\mathrm{DS}_{T}(E^{T})$. ∎
Omitted details from 5.2
In what follows, we provide supporting details for the derivation in 5.2. We recursively define the set of sophisticated dynamically stable matchings for the economy in the example, starting from the case in which all $t=1$ agents wait to be matched. In this example, sophisticated dynamic stability coincides with the rational expectations solution; in particular, the recursion to build the conjectures $\varphi_{\mathrm{SDS}}$, described in the main text, is not necessary.
Note that if nobody matches in $t=1$, a unique stable matching exists in $t=2$, which leads to the matching $m_{\emptyset}$ in the main text. It follows that $m_{\emptyset}$ is the unique sophisticated dynamically stable matching in any two-period economy in which $b_{1}$ is part of the $t=2$-arrivals.
Consider now the case in which two agents in $A_{1}$ wait to be matched, so that the induced economies are $E^{2}_{\setminus\{a_{i},a_{j}\}}$ for $i,j\in\{1,2,3\}$, $i\neq j$. Using the fact that in each of these economies, the conjectures of a period-1 agent who waits to be matched coincides with $m_{\emptyset}$, we conclude the following:
$\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{2},a_{3}\}})$ =$=\{m^{R},m^{C}\},$ $\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{1},a_{3}\}})$ =$=\{m^{L}\},$ $\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{1},a_{2}\}})$ =$=\{m_{\emptyset}\}.$ (B.5)
In $E^{2}_{\setminus\{a_{2},a_{3}\}}$, $a_{1}$ and $b_{1}$ match with each other in $t=2$ if either of them waits to be matched. Because they both discount the future, neither of them blocks $m^{R}$ or $m^{C}$. Instead, $m_{\emptyset}$ is not sophisticated dynamically stable, because $a_{1},b_{1}$ block by matching together in $t=1$. Similarly, in $E^{2}_{\setminus\{a_{1},a_{3}\}}$, agent $a_{2}$ expects to match with $b_{3}$ if they wait to be matched. Both $a_{2},b_{1}$ prefer to match together to waiting to be matched, so that $m^{L}$ is the unique sophisticated dynamically stable matching in that economy. Finally, in $E^{2}_{\setminus\{a_{1},a_{2}\}}$, $a_{3}$ expects to match with $b_{2}$ if they wait to be matched, and this matching is preferred to matching with agent $b_{1}$. Thus, $m_{\emptyset}$ is the unique sophisticated dynamically stable matching.
Consider now the case in which one agent in $A_{1}$ waits to be matched, so that the induced economies are $E^{2}_{\setminus\{a_{i}\}}$ for $i\in\{1,2,3\}$. Using the sophisticated dynamically stable matchings in Equation B.5, we can easily show
$\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{1}\}})$ =$=\{m^{L}\},$ $\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{2}\}})$ =$=\{m^{C},m^{R}\},$ $\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{3}\}})$ =$=\{m^{L}\}.$ (B.6)
We explain the intuition behind $\mathrm{SDS}_{2}(E^{2}_{\setminus\{a_{2}\}})$, because it is used to show why $m^{L}$ is the unique element of $\mathrm{SDS}_{2}(E^{2})$; similar intuition applies for the others. When $a_{2}$ is treated as a period-2 agent, the induced economies by either $a_{1}$ or $a_{3}$ waiting to be matched are $E^{2}_{\setminus\{a_{1},a_{2}\}}$ and $E^{2}_{\setminus\{a_{2},a_{3}\}}$, respectively. Thus, $a_{1}$ can guarantee at most the payoff from matching with $b_{1}$ in $t=2$, whereas $a_{3}$ can guarantee the payoff from matching with $b_{2}$ in $t=2$. It follows that $a_{3}$ must match in $t=2$, and $a_{1}$ must match with $b_{1}$ in $t=1$.
The above arguments imply $a_{2}$ can guarantee at most the payoff from matching with $b_{3}$ in $t=2$ when the economy is $E^{2}$. This justifies that $a_{2}$ does not block $m^{L}$ by waiting, and thus, it is an element of $\mathrm{SDS}_{2}(E^{2})$. Because $a_{1}$ and $a_{3}$ can guarantee the payoffs from matching with $b_{2}$ and $b_{3}$, respectively, it follows that $m^{L}$ is the unique sophisticated dynamically stable matching in $E^{2}$.